OpenStudy (anonymous):

Express log (base 10) x in terms of the natural logarithm?

OpenStudy (anonymous):

\[\log_b(x)=\frac{\log_a(x)}{\log_a(b)}\] is the change o' base formula

OpenStudy (anonymous):

and so \[\log_{10}(x)=\frac{\ln(x)}{\ln(10)}\]

OpenStudy (anonymous):

so the answer is ln base 10 of x?

OpenStudy (anonymous):

\[\ln _{10} (x)\]

OpenStudy (anonymous):

@satellite73 ?

OpenStudy (anonymous):

@artofspeed ln has a base of e

OpenStudy (anonymous):

so what is the answer :S